Joe flew 300 miles with the wind in two hours. after flying agianst the wind for two hours, he had made 270 miles of the return trip. Find thw wind speed and the speed of the plane in still air. Please help me thank you so much!
Answer by ankor@dixie-net.com(15747)

Joe flew 300 miles with the wind in two hours.
After flying against the wind for two hours, he had made 270 miles of the return trip.
Find the wind speed and the speed of the plane in still air.
:
Let s = the speed in still air
Let w = the speed of the wind
then
(s-w) = effective speed against the wind
and
(s+w) = effective speed with the wind
:
write a distance equation for each way (dist = time * effective speed)
2(s+w) = 300
2(s-w) = 270
:
Simplify both equations by dividing both sides by 2, results
s + w = 150
s - w = 135
--------------adding eliminates w, find s
2s = 285
s = 285/2
s = 142.5 mph, plane speed in still air
:
Find w
142.5 + w = 150
w = 150 - 142.5
w = 7.5 mph is the speed of the wind
:
:
YOu can check these solution yourself in the original equations